Fermi National Laboratory
 

Accelerator Update for August 1, 2007 - August 3, 2007

 

Wednesday August 1

The day shift began with Operations monitoring the Tevatron (TeV) store 5589 (56.28E30) and the Recycler stash (217.3E10), with the Antiproton Source (Pbar) stacking (34.9E10), with Linac, Booster, and Main Injector (MI) operating normally, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with MiniBooNE taking beam. NuMI will remain off until after the summer shutdown.

Operations established beam to the Meson MTest experiment at 9:26 AM.

At 9:42 AM, two Booster RF stations (BRF1 & 3) tripped off. They both reset.

Operations reported at 10:04 AM that experts were attempting a fix of the CDF SVX beam monitor high voltage supply without causing a TeV beam abort.

At 10:34 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.

CDF experts reported at 11:31 AM that the SVX problem was with a bad channel in a voltage divider, which they replaced.

Operations reported at 1:13 PM that they put the Meson MTest experiment into its 4 GeV running mode.

At 2:07 PM, Operations reported that a Pbar Debuncher RF station (DRF1-4) had tripped off. An expert investigated and had the station up and running 15 minutes later.

At 3:24 PM, BRF18 tripped off. It reset.

Operations reported at 3:55 PM that the Recycler had stashed antiprotons three times during the day shift.

A Switchyard (SY) expert put MTest into its 2 GeV running mode at 4:56 PM.

MTest finished using the beam for the night at 7:19 PM.

Booster experts began a study period at 9:03 PM. At 10:05 PM, they began a period of no-beam studies.

At 10:10 PM, the Linac Buncher tripped off. It reset.

Booster resumed sending beam at 11:45 PM.

Operations reported at 11:59 PM that the Recycler had stashed two times during the evening shift.

Thursday August 2

The midnight shift began with Operations monitoring store 5589 (31.5E30) and the stash (369.8E10), with Pbar stacking (35.2E10), and with MiniBooNE taking beam.

At 2:35 AM, a MI corrector magnet power supply (I:HT828) tripped off. Operations contacted an expert who instructed operators to replace the chassis.

At 3:04 AM, the MI power supply interlock loop tripped off. It reset.

At 4:08 AM, operators reported that they had replaced the HT828 chassis, but were still unable to run beam. An expert instructed operators on how to run beam without this supply. Operations established beam through MI at 4:20 AM.

At 4:24 AM, BRF11 tripped off. It reset.

MTest resumed taking beam at 6:17 AM.

Operations reported at 7:49 AM that the Recycler had stashed two times during the midnight shift.

Operations terminated store 5589 at 8 AM.

At 8:09 AM, the 750keV line buncher tripped off. It reset.

Operations put the TeV into shot setup at 8:12 AM.

At 8:44 AM, Linac experts reported that LRF3 had tripped off on a LCW leak indication. It reset.

Operations prepared the Meson M02 enclosure for an access at 9:02 AM to fix a broken SWIC.

At 9:51 AM, a TeV RF station (TRF7) tripped off. RF experts responded and had the station back on within 15 minutes.

At 10:45 AM, BRF1 tripped off. It reset.

Operations established store 5590 at 11:11 AM with an initial luminosity of 282.05E30.

At 11:49 AM, BRF11 and the 750 KeV Line Buncher tripped off. They both reset.

Booster experts began a rep rate study at 11:58 AM.

At 1:59 PM, the MiniBooNE horn tripped off for the first of five times in three hours. It reset each time. Operations contacted an expert.

At 2:29 PM, BRF14 tripped off. It reset.

The Neutron Therapy Facility reported at 3:23 PM that they had finished treating their last patient for the day.

At 3:29 PM, BRF16 tripped off. It reset.

At 3:47 PM, BRF1 tripped off. It reset.

MTest completed their use of beam for the day at 6:32 PM.

Operations reported at 11:58 PM that the Recycler had stashed two times during the evening shift.

Friday August 3

The midnight shift began with Operations monitoring store 5590 (78.7E30) and the stash (90.7E10), with Pbar stacking (43.5E10), and with MiniBooNE taking beam.

At 5:28 AM, a Booster vacuum pump tripped off and brought down BRF7 & 8. Everything reset.

MTest resumed taking beam at 6:16 AM.

Operations reported at 8:04 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to stack, stash, and store. The Run Coordinator will allow necessary studies. The last store before the shutdown will be terminated on Saturday evening (8/4/07). TeV experts will conduct studies on Saturday evening that should be completed around midnight. Access to most enclosures will begin around 7 AM on Monday, August 6, 2007. The three main jobs driving this shutdown are the new ComEd power pole installations, the Booster corrector magnet installations, and the MI collimators.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/03/2007    email Fermilab